Phantomdex:

Art galleries sell paintings for a wide range of prices, based on popularity of the artwork. Data collected on the sales of paintings were used to calculate the model ŷ = 1.18(1.027)x, where x represents the age of the painting in years and ŷ represents its predicted selling price in thousands of dollars. Predict the selling price for a work that was painted 80 years ago. $9,943 $47,446 $4,744,561 $9,943,001

Phantomdex:

@midnight97

Midnight97:

Alright on it

Midnight97:

So they've given us our model and our first step is to plug in the age of the painting y=1.18(1.027)^80

Phantomdex:

Alright...makes sense so far.

Midnight97:

After this I am going to have to clean my dogs

crimsonkimchi:

I don't think there's an exponent there. I just see the multiplication. In which case the problem is as simple as multiplying across

Midnight97:

Y=1.18(1.027)^80 =9.943001275

crimsonkimchi:

But that answer makes more sense than multiplying so let's go with that

Midnight97:

To convert from thousands of dollars to just dollars, you'll multiply that by 1000

Midnight97:

Which gives us 9943.001275

Phantomdex:

So it would be D then, since that's the closest.

crimsonkimchi:

It's A, that's the closest

Midnight97:

Let me work please

Phantomdex:

Sorry

Midnight97:

But it is A

Phantomdex:

....yeah that makes more sense that it'd be the smaller number-

Midnight97:

Oi my head

Phantomdex:

Onto the next question

Midnight97:

@phantomdex wrote:
Onto the next question
I have to clean my dogs

Phantomdex:

...crying

Midnight97:

And give my head a bit

Phantomdex:

I don't really have a bit, so is there anyone else online that's good at math that can help me?

crimsonkimchi:

I would hope I'm decent at math

Phantomdex:

Statistics specifically?

Midnight97:

You know what, let's keep on

Midnight97:

My head can wait

Phantomdex:

@midnight97 wrote:
You know what, let's keep on
Are you sure? I don't want you to push yourself for my sake.

crimsonkimchi:

Chances are that I can at least help, I did a lot of math for my undergrad in physics

Midnight97:

@phantomdex wrote:
@midnight97 wrote:
You know what, let's keep on
Are you sure? I don't want you to push yourself for my sake.
Yea

Phantomdex:

Alright if you're sure